# High-Fidelity And Complex Test Data Generation For Real-World SQL Code Generation Services Industry Application & Experience Paper

### v1

@@ 1. Introduction
In the fast-paced landscape of modern data systems, particularly within SQL development and testing, the demand for high-quality mock data that accurately mirrors real-world scenarios is more critical than ever (Whiting et al., 2008; Gartner, 2024; Bhangdiya et al., 2015; Chandra et al., 2015; Somwase et al., 2024; Arcuri, 2019). This need is amplified in industrial settings where data is often confidential and/or sensitive. In such environments, direct access to production data is severely restricted, and traditional approaches like production data sampling become time-consuming, infeasible, or violate data privacy requirements. Historically, mock data generation has relied on manual creation, production data sampling or mathematical models (Somwase et al., 2024; Gartner, 2024; Fögen and Lichter, 2020; Soltana et al., 2017).
@@
One key challenge for effective test data generation lies in achieving realism and diversity within industrial contexts. Mock data must be sufficiently realistic and diverse to effectively simulate the wide range of scenarios encountered in real-world production systems, including both common and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions (Mannino and Abouzied, 2019; Hackolade, [n.d.]). Another challenge is supporting test cases with complex data structure, consistency, and data integrity requirements (Somwase et al., 2024; Chandra et al., 2015; He et al., 2024; Bhangdiya et al., 2015). For industrial applications heavily reliant on structured data types such as JSON, XML, protocol buffers (a.k.a. proto (Contributors, [n.d.])), SQL and Open API schemas, it is paramount that the generated mock data strictly adheres to the specified structure and maintains data integrity and consistency (Endres et al., 2022; Hernandez et al., 2022). Specifically, our SQL code generation services utlizie GoogleSQL (Google Cloud, [n.d.]) that supports complex nested data types, like array, struct and proto. Manual processes are highly susceptible to errors that can compromise the validity of testing and development. Furthermore, scalability and efficiency of generation are critical for enterprise needs, especially when a large amount of test data is needed (Gretel.ai, 2024; Mockaroo, [n.d.]; Hackolade, [n.d.]). As enterprise software systems grow in scale and complexity, the need for large volumes of mock data increases exponentially. Manual and heuristic data generation approaches become impractical and unsustainable in such environments, hindering development velocity and quality assurance efforts.
@@
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as schemas and protocol buffers. The use of LLMs for synthetic data generation is becoming popular in various domains (Singh and Bedathur, 2024; Gretel.ai, 2024; Lu et al., 2023; Karmarkar et al., 2024; Liu et al., 2023a; Zhu et al., 2025; Li et al., 2023; Patel et al., 2024), the generation of structured data introduces unique challenges that are particularly relevant to industrial applications. LLMs, when appropriately trained, can provide realistic and privacy-preserving data synthesis. They can generate mock data that more closely approximates real-world textual data distributions and patterns, capturing subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods (Xu et al., 2024; Sui et al., 2024). This is crucial for realistic testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation (Singh and Bedathur, 2024; Borisov et al., 2022; Sui et al., 2024; Nguyen et al., 2024). Given a schema or proto definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. However, tabular data generation via language models remains an active and evolving field, facing challenges such as accurately capturing complex inter-column (pairwise) relationships, mitigating the risk of “hallucinations” or data inconsistencies, and effectively handling rare values and outliers (also see Table 2 for vanilla LLM evaluation results) (Sui et al., 2024; Nguyen et al., 2024).
@@
Deep generative models (Zhang et al., 2023; Lee et al., 2023; Liu et al., 2023b) have shown promise in tabular data generation, particularly for categorical, numerical, and text attributes. However, they are fundamentally limited by their focus on flat schemas, neglecting nested structures, and their reliance on training data, rendering them infeasible in scenarios where production data is restricted. In contrast, our LLM-based approach, leveraging gemini (AI, [n.d.]) out-of-the-box, directly addresses these limitations by generating structurally compliant data for arbitrarily complex, nested schemas without requiring fine-tuning.
@@
Furthermore, ensuring the semantic validity of the generated data is another crucial challenge that must be addressed to generate truly realistic test data. This is even more critical for testing generated SQL queries (e.g., in NL2SQL systems) on user data. While LLMs might generate data that is syntactically correct and structurally valid, it often lacks the necessary semantic coherence across columns and tables. Most test data generation approaches focus on approximating per-column or pairwise column distributions from the production data (Zhang et al., 2023; Lee et al., 2023; Liu et al., 2023a; Mannino and Abouzied, 2019; Whiting et al., 2008; Bhangdiya et al., 2015; Chandra et al., 2015; Somwase et al., 2024) leading to illogical combinations of column values, semantically incoherent join tables – violating real-world business rules. Overcoming these limitations, including both structural and semantic challenges, is crucial for fully realizing the potential of LLMs in generating comprehensive and accurate mock data for complex workloads.
@@
  • Enhanced Test Coverage for Complex SQL Queries: We demonstrate that our generated test data significantly improves test coverage of Google SQL code generation services, where many test cases are disabled due to limited access to production data and the difficulties of manual data generation for large, complex schemas.
@@
2. SQL code generation services
In this section, we examine the challenges of generating meaningful test data for two real-world GoogleSQL (Google Cloud, [n.d.]) SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
Table 1. Traditional test data generation methods for real-world SQL code generation services. Test coverage is largely limited by the test data quality, especially when the underlying data structure/schema is large and complex with deep nesting.
@@ 2. SQL code generation services
Data generation method Manual Sampling***restricted to a limited set of domains/tables in production
@@
Table 1 summarizes test & schema complexities, as well as the traditional test data generation methods used for the services. The underlying data schema for both services are complex with many columns/fields with nested fields (nested proto and array types). The test complexity for NL2SQL is even more complex as the service generates SQL queries spanning over multiple tables across many customer domains. Test cases are submitted by the internal developers and also by crowdsourcing. The test SQL queries for SQL Code Assistant are mostly simple (single-table) selection queries; however, the manually curated test data overlooks complex nested structures from the schema limiting the test cases to SQL queries targeting mostly flat schemas with non-nested fields. The SQL query complexity vary, but most of them use nested fields and subqueries with joins. SQL code generation services require high-fidelity test data that adheres to the complex schema with nested structures that support complex and diverse sets of SQL queries/functions. Additionally, NL2SQL service evaluation requires testing if predicted SQL generates a result set that is semantically coherent with the natural language question. This also means that the test data instances should be realistic and aligned with the context of the posed data question. Unfortunately, production sampling, though anonymized, is restricted to a few accessible domains and tables, resulting in incomplete test scenarios and low test coverage.
Table 2. Evaluation of test data generation for a large Google Ads production customer table. The table contains 72 columns, 65 of which are nested fields (proto type). Due to data privacy restrictions, production data access and sampling were prohibited. Manual and vanilla LLM generation methods struggled with this moderately complex schema. Our proposed method achieved near-perfect results across all evaluation criteria, demonstrating its ability to generate high-quality test data for wide schemas.
@@
Generation time per instance (50p) 1 hourAn engineer, familiar with the table schema, was tasked with generating a test data instance within one hour. While individual speed may vary, it is practically impossible to curate a comprehensive instance in such a limited timeframe. 37 sec 167 sec
2.1. SQL Code Assistant
@@
Figure 1. Simple selection query accessing deeply nested fields on a fake table, currency and amount.
@@ 2.2. NL2SQL
1Question: Which fake_devices have tasks related to fake_task_name with a start date after 2020-06-01, and what are the tasks and their start dates for the fake_devices? 2-------------------------------------------------------- 3 4SELECT 5 name, 6 ARRAY_AGG(STRUCT(task, start_date)) 7FROM ( 8 SELECT 9 d.name, 10 t.name AS task, 11 FORMAT_TIMESTAMP("%Y-%m-%d", PARSE_TIMESTAMP("%Y-%m-%dT%H:%M:%E*S%Ez", t.start_date)) AS start_date 12 FROM fake_project_tasks AS t 13 INNER JOIN fake_devices AS d 14 ON t.project_id = d.id 15 WHERE ( 16 t.name LIKE ’%fake_task_name’ 17 ) 18 AND t.start_date > "2020-06-01" 19 ORDER BY start_date 20) 21GROUP BY name;
Figure 2. A natural language question and the golden SQL query with a join within the subquery.
@@ 2.3. Problem Definition
Let $\Lambda$ be a schema definition, like table schema or proto, with various types of fields/columns $\Xi$. The column set $\Xi$ includes primitive type (e.g., numerical, string, boolean) $\xi^{prim}$, categorical type $\xi^{cat}$ (e.g., enum), and nested type (e.g., struct, array, proto) $\xi^{nest}$ $\xi^{nest}$ can be another (nested) schema definition $\Lambda_{i}$ that can have more nested structures within $\xi^{nest}_{i}$. Our work focuses on generating arbitrarily complex test data with any number of columns and with deeply nested structures, $\Xi=\{\xi^{prim}_{1},\xi^{cat}_{2},\xi^{nest}_{3},...,\xi_{n}\}$, for testing test SQL queries $\Omega=\{\omega_{1},\omega_{2},...,\omega_{m}\}$ and also without any training tabular dataset $\tau_{\Lambda}$ that confirms to $\Lambda$ – restricted production data access in real-world SQL code generation services.
@@
Specifically, we leverage a pre-trained LLM, gemini, to generate high-fidelity and complex $\tau_{\Lambda}$ that enables testing $\Omega$. Table 2 illustrates the evaluation of test data generation for a large complex schema. Using the out-of-the-box gemini model (Vanilla LLM) for mock test data generation with $\Lambda$ resulted in incomplete data. Our generation method, incorporating strategic pre- and post-processing steps, results in the most complete test data, better than the manually prepared one by a human expert in an hour. Production sampling is not used, since the table is restricted. While the traditional manual data generation is very time-consuming, the LLM based generation runs much faster. We further discuss the per-query test data generation time and cost in Section 6.1.
3. Test data Generation for SQL code generation services
@@
Figure 3. The proposed LLM (gemini) based test data generation process. User provides context map (questions, schema, additional criteria) and test body (SQL query) input and the generator returns the processed generated data that matches criteria.
@@
To automate the generation of test data that aligns with user-specified criteria, we propose a novel test data generation pipeline leveraging the power of LLMs (gemini). Figure 3 provides a high-level overview of this system, illustrating the flow of data and processing from user input to the final validated mock test data output. The process begins with the user providing a context map – containing questions, schema definitions, and additional criteria—along with a test body, such as an SQL query or SQL function implementation. This input triggers a series of modules within the data generation library, including a pre-processor, context extractor, prompt builder, post-processor, and data validator. Each module plays a crucial role in refining and validating the generated data, ensuring it meets both the realized criteria from the schema and the query, as well as the user provided criteria.
3.1. Pre-processor
@@
Figure 4. Pre-processor parses out subqueries (middle) from the originanl SQL query (top) for seeding and independent and dependent data generation. Join information is extracted to highlight the table and column association for the join (bottom).
@@ 3.2. Context extractor
Minimizing edge cases with constraints: We pass constraints extracted directly from the test SQL query, as illustrated in Figure 4, so that the model can explore and generate different values a relevant column in the predicates should take for the test (for instance, the LLM understands that it should generate for currency {``$USD",``GBP",``EUR"\}$ from the query in Figure 1). The same set of constraints are later used for the semantic quality verification of the generated data as well. In Section 6.2 we discuss another example where our LLM-based generation successfully generate test data around $date\_trunc(logdate,quarter)$ constraint, leaving no quarter without test data.
3.3. Prompt builder
@@
(1) Generate the proto column name only if present in the proto file. (2) If you have already generated data for a field, please move to the next field. (3) Generate non-zero values for every numeric field. ”””
@@
Figure 5. Data generation prompt template with system and general instructions. Custom generation seeding instructions can be passed as part of data_generation_signals and user_instruction.
When we make the LLM call for generation, we make the request for each nested field separately on a separate thread (see Section 6.1 for more details). For the primitive/scalar fields, we group and pass one generation call to handle 30 fields at a time. This sometimes leads to unexpected hallucinations, especially when the schema is small with a fewer number of primitive fields (see Section LABEL:sec:error_analysis for an example). To guard against such hallucination, we also keep track of the requested column generations so that any hallucinated columns can be detected and removed.
3.4. Post-processor
This module can add the data constraints and variations by copying the basic data generated by the LLM or existing data written by the user. Since the constraints and variations will affect only a smaller number of columns in a proto, we could easily just modify those components.

### v2

@@ I Introduction
In the fast-paced landscape of modern data systems, particularly within SQL development and testing, the demand for high-quality mock data that accurately mirrors real-world scenarios is more critical than ever [1, 2, 3, 4, 5, 6]. This need is amplified in industrial settings where data is often confidential and/or sensitive. In such environments, direct access to production data is severely restricted, and traditional approaches like production data sampling become time-consuming, infeasible, or violate data privacy requirements. Historically, mock data generation has relied on manual creation, production data sampling or mathematical models [5, 2, 7, 8].
@@
One key challenge for effective test data generation lies in achieving realism and diversity within industrial contexts. Mock data must be sufficiently realistic and diverse to effectively simulate the wide range of scenarios encountered in real-world production systems, including both common and rare edge cases. Existing data generation methods often fall short in capturing this diversity, while rule-based systems tend to be inflexible and limited in their ability to generate varied data that reflects the complexities of industrial data distributions [9, 10]. Another challenge is supporting test cases with complex data structure, consistency, and data integrity requirements [5, 4, 11, 3]. For industrial applications heavily reliant on structured data types such as JSON, XML, protocol buffers (a.k.a. proto [12]), SQL and Open API schemas, it is paramount that the generated mock data strictly adheres to the specified structure and maintains data integrity and consistency [13, 14]. Specifically, our SQL code generation services utlizie GoogleSQL [15] that supports complex nested data types, like array, struct and proto – making it a representative case for the challenges faced by many modern SQL dialects that support nested data structures. Manual processes are highly susceptible to errors that can compromise the validity of testing and development. Furthermore, scalability and efficiency of generation are critical for enterprise needs, especially when a large amount of test data is needed [16, 17, 10]. As enterprise software systems grow in scale and complexity, the need for large volumes of mock data increases exponentially. Manual and heuristic data generation approaches become impractical and unsustainable in such environments, hindering development velocity and quality assurance efforts.
@@
The emergence of Large Language Models (LLMs) presents a paradigm shift, offering new avenues for automating the generation of mock data, especially for structured formats such as schemas and protocol buffers. The use of LLMs for synthetic data generation is becoming popular in various domains [18, 16, 19, 20, 21, 22, 23, 24], the generation of structured data introduces unique challenges that are particularly relevant to industrial applications. LLMs, when appropriately trained, can provide realistic and privacy-preserving data synthesis. They can generate mock data that more closely approximates real-world textual data distributions and patterns, capturing subtle nuances and semantic relationships between data attributes that are difficult to replicate with traditional methods [25, 26]. This is crucial for realistic testing while also avoiding the privacy risks associated with using actual production data (regulatory security and privacy policy restricts access to user data). Additionally, LLMs can understand SQL structures and facilitate structurally compliant and schema-aware data generation [18, 27, 26, 28]. Given a schema or proto definition, LLMs can be guided to generate mock data that adheres to the specified structure and constraints. However, tabular data generation via language models remains an active and evolving field, facing challenges such as accurately capturing complex inter-column (pairwise) relationships, mitigating the risk of “hallucinations” or data inconsistencies, and effectively handling rare values and outliers (also see Table II for vanilla LLM evaluation results) [26, 28].
@@
Deep generative models [29, 30, 31] have shown promise in tabular data generation, particularly for categorical, numerical, and text attributes. However, they are fundamentally limited by their focus on flat schemas, neglecting nested structures, and their reliance on training data, rendering them infeasible in scenarios where production data is restricted. In contrast, our LLM-based approach, leveraging gemini [32] out-of-the-box, directly addresses these limitations by generating structurally compliant data for arbitrarily complex, nested schemas without requiring fine-tuning.
@@
Furthermore, ensuring the semantic validity of the generated data is another crucial challenge that must be addressed to generate truly realistic test data. This is even more critical for testing generated SQL queries (e.g., in NL2SQL systems) on user data. While LLMs might generate data that is syntactically correct and structurally valid, it often lacks the necessary semantic coherence across columns and tables. Most test data generation approaches focus on approximating per-column or pairwise column distributions from the production data [29, 30, 21, 9, 1, 3, 4, 5] leading to illogical combinations of column values, semantically incoherent join tables – violating real-world business rules. Overcoming these limitations, including both structural and semantic challenges, is crucial for fully realizing the potential of LLMs in generating comprehensive and accurate mock data for complex workloads.
@@
  • Enhanced Test Coverage for Complex SQL Queries: We demonstrate that our generated test data significantly improves test coverage of Google SQL code generation services, where many test cases are disabled due to limited access to production data and the difficulties of manual data generation for large, complex schemas. Our fundamental approach is dialect-agnostic and should be readily applicable to other SQL code generation services that face similar challenges.
@@
II SQL code generation services
In this section, we examine the challenges of generating meaningful test data for two real-world GoogleSQL [15] SQL code generation services: SQL Code Assistant and NL2SQL. These services represent distinct use cases with varying complexities, highlighting the limitations of traditional test data generation methods which produces data that yields low coverage of the GoogleSQL queries under test.
TABLE I: Traditional test data generation methods for real-world SQL code generation services. Test coverage is largely limited by the test data quality, especially when the underlying data structure/schema is large and complex with deep nesting.
@@ II SQL code generation services
Data generation method Manual Sampling111restricted to a limited set of domains/tables in production
@@
Table I summarizes test & schema complexities, as well as the traditional test data generation methods used for the services. The underlying data schema for both services are complex with many columns/fields with nested fields (nested proto and array types). The test complexity for NL2SQL is even more complex as the service generates SQL queries spanning over multiple tables across many customer domains. Test cases are submitted by the internal developers and also by crowdsourcing. The test SQL queries for SQL Code Assistant are mostly simple (single-table) selection queries; however, the manually curated test data overlooks complex nested structures from the schema limiting the test cases to SQL queries targeting mostly flat schemas with non-nested fields. The SQL query complexity vary, but most of them use nested fields and subqueries with joins. SQL code generation services require high-fidelity test data that adheres to the complex schema with nested structures that support complex and diverse sets of SQL queries/functions. Additionally, NL2SQL service evaluation requires testing if predicted SQL generates a result set that is semantically coherent with the natural language question. This also means that the test data instances should be realistic and aligned with the context of the posed data question. Unfortunately, production sampling, though anonymized, is restricted to a few accessible domains and tables, resulting in incomplete test scenarios and low test coverage.
TABLE II: Evaluation of test data generation for a large Google Ads production customer table. The table contains 72 columns, 65 of which are nested fields (proto type). Due to data privacy restrictions, production data access and sampling were prohibited. Manual and vanilla LLM generation methods struggled with this moderately complex schema. Our proposed method achieved near-perfect results across all evaluation criteria, demonstrating its ability to generate high-quality test data for wide schemas.
@@
Generation time per instance (50p) 1 hour [^2] An engineer, familiar with the table schema, was tasked with generating a test data instance within one hour. While individual speed may vary, it is practically impossible to curate a comprehensive instance in such a limited timeframe. 37 sec 167 sec
II-A SQL Code Assistant
@@
Figure 1: Simple selection query accessing deeply nested fields on a fake table, currency and amount.
@@ II-B NL2SQL
1Question: Which fake_devices have tasks related to fake_task_name with a start date after 2020-06-01, and what are the tasks and their start dates for the fake_devices? 2---------------------------------------------------- 3 4SELECT 5 name, 6 ARRAY_AGG(STRUCT(task, start_date)) 7FROM ( 8 SELECT 9 d.name, 10 t.name AS task, 11 FORMAT_TIMESTAMP("%Y-%m-%d", PARSE_TIMESTAMP("%Y-%m-%dT%H:%M:%E*S%Ez", t.start_date)) AS start_date 12 FROM fake_project_tasks AS t 13 INNER JOIN fake_devices AS d 14 ON t.project_id = d.id 15 WHERE ( 16 t.name LIKE ’%fake_task_name’ 17 ) 18 AND t.start_date > "2020-06-01" 19 ORDER BY start_date 20) 21GROUP BY name;
Figure 2: A natural language question and the golden SQL query with a join within the subquery.
@@ II-C Problem Definition
Let $\Lambda$ be a schema definition, like table schema or proto, with various types of fields/columns $\Xi$. The column set $\Xi$ includes primitive type (e.g., numerical, string, boolean) $\xi^{prim}$, categorical type $\xi^{cat}$ (e.g., enum), and nested type (e.g., struct, array, proto) $\xi^{nest}$ . $\xi^{nest}$ can be another (nested) schema definition $\Lambda_{i}$ that can have more nested structures within $\xi^{nest}_{i}$. Our work focuses on generating arbitrarily complex test data with any number of columns and with deeply nested structures, $\Xi=\{\xi^{prim}_{1},\xi^{cat}_{2},\xi^{nest}_{3},...,\xi_{n}\}$, for testing test SQL queries $\Omega=\{\omega_{1},\omega_{2},...,\omega_{m}\}$ and also without any training tabular dataset $\tau_{\Lambda}$ that confirms to $\Lambda$ – restricted production data access in real-world SQL code generation services.
@@
Specifically, we leverage a pre-trained LLM, gemini, to generate high-fidelity and complex $\tau_{\Lambda}$ that enables testing $\Omega$. Table II illustrates the evaluation of test data generation for a large complex schema. Using the out-of-the-box gemini model (Vanilla LLM) for mock test data generation with $\Lambda$ resulted in incomplete data. Our generation method, incorporating strategic pre- and post-processing steps, results in the most complete test data, better than the manually prepared one by a human expert in an hour. Production sampling is not used, since the table is restricted. While the traditional manual data generation is very time-consuming, the LLM based generation runs much faster. We further discuss the per-query test data generation time and cost in Section VI-A.
III Test data Generation for SQL code generation services
@@
Figure 3: The proposed LLM (gemini) based test data generation process. User provides context map (questions, schema, additional criteria) and test body (SQL query) as input. This data generation request triggers a series of data processing steps and modules. The goal is to generate syntactically correct and semantically relevant test data for SQL code generation services, like NL2SQL. We leverage the latest LLM capabilities to do this efficiently at scale.
@@
To automate the generation of test data that aligns with user-specified criteria, we propose a novel test data generation pipeline leveraging the power of LLMs (gemini). Figure 3 provides a high-level overview of this process, illustrating the flow of data and processing from user input to the final validated mock test data output. The process begins with the user providing a context map – containing questions, schema definitions, and additional criteria—along with a test body, such as an SQL query or SQL function implementation. This input triggers a series of modules within the data generation library, including a pre-processor, context extractor, prompt builder, post-processor, and data validator. Each module plays a crucial role in refining and validating the generated data, ensuring it meets both the realized criteria from the schema and the query, as well as the user provided criteria.
III-A Pre-processor
@@
Figure 4: Pre-processor parses out subqueries (middle) from the original SQL query (top) for seeding and independent and dependent data generation. Join information is extracted to highlight the table and column association for the join (bottom).
@@ III-B Context extractor
Minimizing edge cases with constraints: We pass constraints extracted directly from the test SQL query, as illustrated in Figure 4, so that the model can explore and generate different values a relevant column in the predicates should take for the test (for instance, the LLM understands that it should generate for currency $\in\{``USD",``GBP",``EUR$"\} from the query in Figure 1). The same set of constraints are later used for the semantic quality verification of the generated data as well. In Section VI-B we discuss another example where our LLM-based generation successfully generate test data around date_trunc(logdate,quarter)date\_trunc(logdate,quarter) constraint, leaving no quarter without test data.
III-C Prompt builder
@@
1. Generate the proto column name only if present in the proto file. 2. If you have already generated data for a field, please move to the next field. 3. Generate non-zero values for every numeric field. ”””
@@
Figure 5: Data generation prompt template with system and general instructions. Custom generation seeding instructions can be passed as part of data_generation_signals and user_instruction.
1---------------------------------------------------- 2# Without column grouping: 3 4json[{ 5 "contract_created_date": "2023-08-15", 6 "contract_created_month": "2023-08" 7 "contract_created_year": 2023, 8 "contract_effective_date": "2023-04-01", 9 "contract_closed_date": "2023-03-30", 10 "contract_executed_date": "2023-03-29", 11 "contract_expiry_date": "2024-03-31", 12 "fully_signed_time": "2023-03-29T10:00:00Z", "contract_termination_date": "2023-12-31", 13}, 14...] 15---------------------------------------------------- 16# With column grouping: 17 18json[{ 19 "contract_created_date": "2023-10-27", 20 "contract_created_month": 10, 21 "contract_created_year": 2023, 22 "contract_effective_date": "2023-11-01", 23 "contract_closed_date": "2024-10-31", 24 "contract_executed_date": "2023-10-30", 25 "contract_expiry_date": "2025-10-31", 26 "fully_signed_time": "2023-10-30T10:00:00Z", 27 "contract_termination_date": "2024-05-15", 28 "contract_termination_date_option": "SpecificDate", 29 "contract_months_from_effective_date": 12, 30 "simba_contract_create_date": "2023-10-27", 31}, 32...]
Figure 6: Column grouping based on column correlation helps generating more coherent set of values. Without column grouping (above) the contract effective (start) date is before the contract created date; with column grouping (below) all dates are aligned and correct.
1---------------------------------------------------- 2# column annotation 3 4"terminated_by_clm_system" // "Ifthecontractwasterminated,thisfieldindicatestheContractLifecycleManagementSystemofthecontract(s)responsibleforitstermination.Thisfieldispartofthesourceinformationaboutthiscontract’slineagewithintheContractFamilyLinkage." 5 6---------------------------------------------------- 7# Generated column value without annotation: 8"terminated_by_clm_system": "TerminationSystem", 9 10---------------------------------------------------- 11# Generated column value with annotation: 12"terminated_by_clm_system": "InternalCLMSystemCode",
Figure 7: An example column annotation prepared by the context extractor from human written comments, source code and documentation. This allows the generation to look up and consider valid set of column values, instead of random (hallucinated) strings.
When we make the LLM call for generation, we make the request for each nested field separately on a separate thread (see Section VI-A for more details). For the primitive/scalar fields, we group and pass one generation call to handle 30 fields at a time. The grouping is based on the column correlations identified by the context extractor. This helps maintain semantic consistency across the correlated columns, as supposed to generating a valid value for each column individually (see Figure 6). The column grouping sometimes leads to unexpected hallucinations where the LLM introduces fake (non-existent) columns, especially when the schema is small with a fewer number of primitive fields (see Section VI-C for an example). To guard against such hallucination, we also keep track of the requested column generations so that any hallucinated columns can be detected and removed. The schema (proto description) passed here also contains column/field annotations prepared by the context extractor. The annotations are attached to the corresponding schema elements as comments. This also improves the semantics of the generated data, as illustrated in Figure 7.
III-D Post-processor
This module can add the data constraints and variations by copying the basic data generated by the LLM or existing data written by the user. The extracted join constraints in the pre-pro in the pre-processing are also enforced by the post-processor, where the column values in the joined tables at the join columns are matched and copied from one table data to another. Since the constraints and variations will affect only a smaller number of columns in a proto, we could easily just modify those components.